OpenStudy (anonymous):

Factor completely: 2x3 + 14x2 + 4x + 28

OpenStudy (anonymous):

first 2 is a factor: = 2(x^3 + 7x^2 + 2x + 14) use factor theorem :- this is a tedious one f(-7) = 0 so (x+ 7) is a factor

OpenStudy (deoxna):

First, factor out the 2: 2(x^3+7x^2+2x+14) Then, divide the polynomial in the parentheses by one of its factors (x+7), which you can find by plugging in one of the factors of the last number of the polynomial until you get 0. So when you plug -7 into the polynomial, you get 0. So then divide it by x+ the opposite of 7 which is +7 (x+7): ____________________ x+7|x^3+7x^2+2x+14 = x^2+2 So: x^3+7x^2+2x+14 = (x^2+2)(x+7) And that factors to: 2(x+2)(x+2)(x+7)